Jim and Al sold a total of 600 hotodgs in a month. If Jim sold 5 times as many as Al. How much did each sell?
Found 2 solutions by mananth, iyah98:
Answer by mananth(16946)

let Al sell x hotdogs
Jim sold 5x
...
5x+x = 600
6x=600
/6
x= 100 sold by Al
500 sold by Jim
...
m.ananth@hotmail.ca

Answer by iyah98(2)

equation base on the given statement
J + A = 600
J = 5A
Simplified
5A + A = 600
6A = 600
A = 100 Hotdogs sold by Al
Jim sold 500 hotdogs